OpenStudy (anonymous):

Determine wheather each functio is linear. If so, give the slope and the y-intercept of the functions graph. PLEASE SHOW WORK 1. 4y=2x 2. 5=8-3y 3. 4y=10-1/x

OpenStudy (campbell_st):

rewrite each in the form y = mx + b so the 1st is y = 1/2 x 2nd is y = 1 3rd is not linear.... now the coeffiecient of x is the slope... and constant is the y-intercept alternatively.... graph the functions and see which are straight lines